find three consecutive number such that six times the smallest number is equal to double the sum of other two.
Answer by addingup(3677)

We have three consecutive numbers, represented by: x, x+1, x+2. Then:
6x= 2(x+1+x+2)
6x= 2(2x+3)
6x= 4x+6
2x= 6
x= 3
The numbers are 3, 4, and 5
Proof:
6*3= 18 and 2(4+5)= 18